0.28g of monochloroethanoic acid CH2ClOOH is dissolved in 250cl of distilled water. The pH of the solution obtained is 2.4.

What is the concentration of the monochloroethanoic acid?

What is the value of the equilibrium constant of the reaction of the monochloroethanoic acid with water?


1
Expert's answer
2021-07-20T11:05:01-0400

Molar Mass of CH2ClOOH = 94.96

0.28 × 94.96 = 26.589

Concentration = 0.28/0.25 = 1.12mol/L